OpenStudy (anonymous):

Write an equation in point slope form given: slope, m, = -1/5 and point (5, -2). A. y + 2 = -(1/5)(x - 5) B. y + 5 = -(1/5)(x - 2) C. y – 5 = (1/5)(x + 2) D. y – 5 = (1/5)(x - 2)

OpenStudy (anonymous):

The general form for point slope is:\[y-y_1 = m(x-x_1)\] Can you use the information given to plug in what you know to this equation?

OpenStudy (anonymous):

i dont know how

OpenStudy (anonymous):

Okay! Lets work on this. You know that m = slope, and you know that your slope is -1/5. So the first thing to do would be to plug in m to your equation to get: \[y - y_1 = -1/5(x-x_1)\] now, the x1 and y1 are the x and y values from the ordered pair you're given. You do the same thing we did with the slope and plug those in for x1 and y1.

OpenStudy (anonymous):

1-y=-1/5(1-x)